OpenStudy (anonymous):

what is the point- slope form of (-2,11) and (6,7). show me that step only because i keep getting it wrong

OpenStudy (anonymous):

slope=m=(7-11)/(6+2) m=-5/8 will be the slope.

OpenStudy (anonymous):

Point-slope form is: \[y - y _{0} = m (x - x _{0})\] Ravi2 showed you how to obtain the slope, m. Now, to have the complete point-slope form, you can pick one of these points as your (x0, y0): (-2,11) or (6,7). The rest just involves you plugging in the (x0,y0) and your slope, m into the point-slope form.